OpenStudy (anonymous):

How to solve this integral-> ∫ [0, x] (1)/(1+t^4) dt=

OpenStudy (luigi0210):

Use the fundamental theorem of calculus, part 1.

OpenStudy (anonymous):

I can't grasp this one the answer includes arctan or ln

OpenStudy (amistre64):

substitute .... let t^2 = x maybe?

OpenStudy (anonymous):

hmm I'll try it

OpenStudy (amistre64):

not =x but something more approriate of course

OpenStudy (anonymous):

you mean [0, t^2]

OpenStudy (amistre64):

\[\int_{0}^{x}\frac{1+t^4-t^4}{1+t^4}dt\] \[\int_{0}^{x}1-\frac{t^4}{1+t^4}dt\] if t^2 = a 2t dt = da, t dt = da/2t \[\int_{0}^{x}1-\frac{t^4}{1+a^2}\frac{da}{2t}\] \[\int_{0}^{x}1-\frac{t^3}{1+a^2}\frac{da}{2}\] \[\int_{0}^{x}1-\frac{a\sqrt a}{1+a^2}\frac{da}{2}\] might be better with a trig sub .... or may just be a pain regardless

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+1%2F%281%2Bx%5E4%29 the wolf says its ugly .... so the question is, do we need to integrate it or is this part of taking the derivative of th eintegration?

OpenStudy (amistre64):

1-t^4+t^8-t^12 ... -------------------- 1 + t^4 | 1 -(1+t^4) -------- -t^4 \[\int(1-t^4+t^8-t^{12}+ ...)\] \[t-\frac 15t^5+\frac19t^9-\frac1{13}t^{13}+ ...\] at t=0 is 0, so let t=x

OpenStudy (anonymous):

Hmmm bit more than I can handle that's for sure

OpenStudy (anonymous):

Is there are way I do a reversal trick with the answers given

OpenStudy (anonymous):

A way I can do*

OpenStudy (amistre64):

im not too sure what you mean by that

OpenStudy (anonymous):

Okay so I have the answer chooses is there an inverse I can to on the choices that would will give me the original integral

OpenStudy (amistre64):

what are the choices ... that might help me see a solution better

OpenStudy (anonymous):

1/2 arctan(x^2) 1/2 arctan(x^4) 1/2 ln(1+x^4) 1/2arcsin(x^2)

OpenStudy (amistre64):

well, i know what the derivative will be .....\[\frac{d}{dx}\left(\int_{0}^{x}f(t)~dt\right)=\frac{d}{dx}\left[F(x)-F(0)\right]=f(x)x'-f(0)0'=f(x)x'\]

OpenStudy (amistre64):

so, which of these answers has a the derivative of x'/(1+x^4) ?

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

so the first one is x/x^4 +1

OpenStudy (anonymous):

the rest of them don't match

OpenStudy (amistre64):

so A is the best bet, unless we can work some algebra on the others perhaps?

OpenStudy (anonymous):

No no I can't so I'll choose a

OpenStudy (anonymous):

Thanks though

OpenStudy (amistre64):

the wolfs answer is ugly, not as pretty as those so its hard to dbl check :)

OpenStudy (anonymous):

Yeah I know I tried other calculators

OpenStudy (amistre64):

the partial fraction is messy too, which then leads to the messy integral lol